choose your preferred solution method and solve this quadratic equation:
6x^2 + 7x = 3

Answer by Alan3354(69443)

6x^2 + 7x = 3
6x^2 + 7x - 3 = 0
(2x + 3)*(3x - 1) = 0
x = -3/2
x = 1/3